how much water should be added to 300 litres of a solution so that its concentration may change from 25% to 20%?
Found 2 solutions by Alan3354, richwmiller:
Answer by Alan3354(69443)

Add W liters of water.
300*25 + W*0 = (300 + W)*20
7500 = 20W + 6000
W = 75 liters

Answer by richwmiller(17219)

from the point of view of water
300*.75+x=(300+x)*(.80)
x=75